Fill in the boxes with the correct symbol out of >,
 -4/5 [ ] -5/7

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The LCM of the denominators 5 and 7 is 35
∴(-4/5) = [(-4 × 7)/ (5 × 7)]= (-28/35)
And (-5/7)= [(-5 × 5)/ (7 × 5)]
= (-25/35)
Now,
-28 < -25
So, (-28/35) < (- 25/35)
Hence, -4/5 [<] -5/7

Thanks